# ETH Zurich

### v1

@@ 2.1 Data sources
2.2 Construction of classification tasks
We construct a set of classification tasks using the variables of the Supreme Court and the U.S. Courts of Appeals databases. We construct a total of 260 distinct classification tasks, 38 of them corresponding to the Supreme Court database and 232 to the U.S. Court of Appeals. For each of these classification tasks, we construct a prompt template consisting of a general description of the task, followed by a multiple choice question containing each of the possible variable codes. We formulate the task description, question, and answer choices by closely following the language of the variable description of the databases’ documentation. Thereafter, for every case contained in these databases, we use the provided case citations to search for its corresponding majority opinion of the court on the Caselaw Access Project , a database of digitized court opinions.
We divide court opinions into a 70%/10%/20% train/validation/test split. Since many of the classification taks contain heavily imbalanced classes, we subsample the majority class such that there are at most as many task examples in the majority class as task examples in all other classes combined. Therefore, a constant classifier that outputs the majority class label will never achieve more than $50\%$ accuracy on any task. We find that on average over all tasks, the majority classifier has 41.7% accuracy. We plot some statistics of the tasks in Figure 2. First, court opinions tend to be long, with 12% having above 8,000 tokens, the typical maximum context size for current state-of-the-art models, such as Llama 3. Second, some tasks have a large number of classes, with 28% of tasks having more than 10 classes. Third, there is a large variability in terms of the number of task examples, ranging from a couple dozen to 18500 task examples. In the remainder of the article, we highlight ten different tasks, six from the SCDB and four from the USCAD. These tasks differ in complexity, ranging from relatively simple tasks (e.g., determining the issue area) to relatively complex ones (e.g., determining the ideological “direction” of the court decision). Section 2.4 contains a brief description of these tasks, Section B the full list of all tasks.
Example of the prompt template.
We use a prompt template identical to the one for the MMLU benchmark (Hendrycks et al., 2020). We provide as an example the prompt template corresponding to the Supreme Court precedent alteration variable (“sc_precedentalteration” in Section B).
What follows is an opinion from the Supreme Court of the United States.
Your task  is to identify  whether the opinion effectively says that
the decision in this case "overruled" one or more of the Court\’s own
precedents. Alteration also extends to language in the majority opinion
that states that a precedent of the Supreme Court has been "disapproved,"
or is "no longer good law". Note, however, that alteration does not
apply to cases in which the Court "distinguishes" a precedent.
@@ 2.2 Construction of classification tasks
 Question: Did the the decision of the court overrule one or more
 of the Court’s own precedents?
@@ A. Yes
 Answer:
Due to diverse set of models and very large number of tasks under consideration, we decided to perform no prompt tuning. [^3] Multiple choice prompts are popular for language model evaluation because they only requires a single forward pass per example. Since many popular benchmarks are phrased as multiple-choice, recent instruction tuned models tend to do well for them. Note that more involved prompting strategies –e.g., chain-of-thought (Wei et al., 2022)– can yield better task performance but are substantially more expensive. For legal tasks specifically, the choice of prompt can have a significant effect in performance (Li et al., 2024). Prompt tuning requires task-specific domain knowledge and can be reasonably time consuming. Instead, we formulate the task description, question, and answer choices by closely following the language of the variable description of the databases’ documentation.
2.3 Zero-shot and few-shot evaluation baselines
Figure 3: Zero-shot accuracy of models on different legal classification tasks. Top: Mean accuracy over all cases. Middle: Mean task accuracy. Bottom: Accuracy in 10 representative tasks. Error bars indicate 90% Clopper-Pearson exact Binomial confidence intervals.
To establish baselines, we evaluate the zero-shot performance of different language models on all 260 classification tasks. Figure 3 presents the accuracies achieved in different task categories (all, Supreme Court, Court of Appeals), as well as ten representative tasks. We find that only Llama 3 70B Inst and GPT-4 are consistently better than the constant classifier that outputs the majority class. Figure 4 compares the zero-shot accuracy of GPT-4 and Llama 3 70B across all 260 tasks.
Figure 4: Difference in zero-shot accuracy between GPT4 and Llama 3 70B Instruct. Each vertical bar represents the accuracy difference on one task, sorted in ascending order.
Confidence intervals.
There are two natural ways to compute aggregate accuracies across multiple tasks. One is to compute the accuracy across all cases in the task collection. Tasks with fewer cases have lower weight under this average. The other is to compute the accuracy on each task separately and average those out. Since some tasks have far fewer documents than others (see Figure 2), the latter has higher variance. We indicate this by plotting the average error bars across tasks. Hence, the error bars for mean task accuracy are larger. The error bars for the case average are generally negligible, while several very small tasks have large error bars.
Table 1: Zero-shot and few-shot accuracies for different models
Model All tasks Supreme Court Court of Appeals
LegalBERT 24.58 13.59 26.45
Saul 7B Instruct 34.41 20.23 36.82
Mistral 7B Instruct 39.90 19.52 43.36
Majority class 41.72 31.48 43.45
Llama 3 8B Instruct 42.58 32.80 44.24
Mixtral 8x7B Instruct 43.22 24.37 46.42
Llama 3 70B Instruct 58.37 47.13 60.27
GPT-4 62.89 59.78 63.42
GPT-4 $32k$ $3$-shot 58.38 53.99 59.12
Multi-shot prompting.
We also evaluated whether multi-shot prompting would lead to improvements. As Figure 2 shows, many documents have thousands of tokens. The default context window for GPT-4 is 8,000 tokens. A version with 32,000 tokens is available at twice the cost. We evaluated the $32k$ version with $3$-shot prompting. This, however, did not yield any improvements over zero-shot prompting the standard model. See the last row in Table 1.
2.4 Highlighted tasks
@@
Throughout this paper, as in Figure 3, we provide detailed results for ten tasks. Six of these tasks are from the SCDB, and four are from the USCAD. We selected tasks that we believe are particularly relevant to the legal community and chose tasks with varying levels of complexity, ranging from relatively simple (e.g., determining the issue area) to more complex (e.g., determining the ideological ’direction’ of the court decision).
Four tasks from the USCAD and all tasks from the SCDB were selected to form pairs, with each pair consisting of one task from the USCAD and one from the SCDB that capture similar concepts. It is important to note that, despite capturing broadly similar concepts, the precise formulation of the tasks might differ between the USCAD and the SCDB, making them less than perfectly comparable. In addition to the four pairs, we include two tasks from the SCDB that involve determining features of the decision reviewed by the Supreme Court on the basis of the Supreme Court opinion.
The following is a description of the task pairs:
3 Fine-tuning and the power of specialization
Figure 5: Zero-shot accuracy of four different models in comparison with the accuracy of (fine-tuned) Lawma 8B and 70B on different legal classification tasks. Top: Mean accuracy over all cases. Middle: Mean task accuracy. Bottom: Accuracy in 10 representative tasks. Error bars indicate 90% Clopper-Pearson exact Binomial confidence intervals.
In this section, we present a detailed analysis of fine-tuning on different tasks. We start by fine-tuning Llama 3 8B Inst and Llama 3 70B Inst on all tasks simultaneously, resulting in our Lawma 8B and Lawma 70B models. We then perform additional fine-tuning experiments highlighting different aspects, including the scaling behaviour of fine-tuning larger base models, the sample efficiency of fine-tuning, its generalization to unseen tasks and Courts, and the effect of single task specialization.
We first fine-tune Llama 3 8B Inst and Llama 3 70B Inst on all tasks simultaneously. We refer to these fine-tuned models as Lawma 8B and Lawma 70B, respectively. We fine-tune on the 260 classification tasks described in Section 2.2. The fine-tuning dataset contains a total of 1.96B tokens. For Lawma 3 8B, we fine-tune for 3 epochs. For Lawma 3 70B, we fine-tune for 1 epoch. We find that additional epochs hurt performance. See Appendix A for additional details regarding the model training.
We compare in Figure 5 the task accuracies of Lawma 8B and 70B to that of their respective base models Llama 3 8B Inst and Llama 3 70B Inst, as well as GPT-4. Fine-tuning leads to large improvements in average task accuracy: Lawma 8B outperforms Llama 3 8B Inst by 37.7 accuracy points and Lawma 70B outperforms Llama 3 70B by 23.5 accuracy points. Both Lawma 8B and Lawma 70B outperform GPT4, Lawma 8B on average by 17.3 accuracy points and Lawma 70B on average by 19.0 accuracy points.
@@ 3 Fine-tuning and the power of specialization
Figure 6: Distribution of task performance across all tasks for Llama 3, GPT4, and Lawma.
Moreover, the Lawma models are the top performing models in each of the 10 highlighted tasks. In fact, the both Lawma 8B and Lawma 70B outperforms GPT4 in about 95% of all tasks, see Figure 1. Figure 6 further demonstrates the large effect of fine-tuning by showing the histogram of task accuracies achieved by Lawma 8B and Lawma 70B in comparison with Llama 3 8B Inst, Llama 3 70B Inst, and GPT4.
We find that Lawma 8B closely matches the performance of Lawma 70B. Specifically, Lawma 70B outperforms Lawma 8B in terms of mean case accuracy by only 0.9 accuracy points for Supreme Court tasks and by 1.2 accuracy points for Appeals Court tasks (Figure 5). In terms of mean task accuracy, the difference between Lawma 8B and Lawma 70B is not statistically significant. We find that by training the 8B model on more epochs than the 70B model, it attains a similar loss as the 70B model. Further training either model results in overfitting and thus performance degrades on average. This suggests that, for our fine-tuning dataset, further scaling model size (e.g., fine-tuning GPT4) is unlikely to yield major improvements. A more promising direction is to instead improve the diversity and quantity of the fine-tuning data. On the flip side, practitioners may choose to use Lawma 8B instead of the 70B model with little cost in model performance.
3.1 Performance after fine-tuning scales with pretraining compute
The performance of specialized models tends to scale with pretraining compute (Dominguez-Olmedo et al., 2024). We investigate how performance after fine-tuning scales with the pretraining compute of the base model. We fine-tune the following models for a single epoch: Pythia 70M, Pythia 160M, Pythia 410M, Pythia 1B, Pythia 2.8B, Pythia 6.9B (Biderman et al., 2023), Llama 2 7B (Touvron et al., 2023), Llama 3 8B Inst and Llama 3 70B Inst. We fine-tune on all 260 tasks simultaneously. We approximate pretraining compute in FLOPs as $C\approx 6\cdot N\cdot D$ (Kaplan et al., 2020), where $N$ is model size and $D$ is the number of pretraining tokens.
Figure 7: Performance after one epoch of fine-tuning increases monotonically in the amount of pretraining of the base model. Models left to right (blue dots): Pythia 70M, Pythia 160M, Pythia 410M, Pythia 1B, Pythia 2.8B, Pythia 6.9B, Llama 2 7B, Llama 3 8B Inst, Llama 3 70B Inst.
We find that mean task accuracy after fine-tuning improves with pretraining compute (Figure 7). However, we find signs of diminishing returns. For the Supreme Court tasks, performance increases steadily from $10^{20}$ to $10^{24}$ FLOPs, but further scaling to $10^{25}$ FLOPs only improves performance by an additional 4.0 accuracy points. For Appeals Court tasks, performance sharply increases from $10^{20}$ to $10^{21}$ FLOPs (i.e., Pythia 1B – which interestingly already beats GPT-4 zero-shot), but stagnates thereafter, only improving by an additional 8.5 accuracy points when scaling to $10^{25}$ FLOPs (i.e., Llama 3 70B Inst).
Our findings suggest that major improvements will likely not come from model scale alone. Rather, future work should focus on obtaining better scaling behavior. One promising direction is to improve the quality, quantity and diversity of the fine-tuning data.
@@ 3.2 Sample efficiency
We study how task accuracy scales as models fine-tune on more training examples. We consider the 10 tasks highlighted in Section 2.4. We fine-tune Llama 3 8B Instruct on each task independently, rather than on all tasks simultaneously as in the previous experiments. For each task, we fine-tune on 10, 50, 100, 250, 500, and 1000 task examples. We select task examples uniformly at random, and train 5 different models corresponding to different random seeds on the examples selected for training. We therefore fine-tune and evaluate a total of $10\cdot 6\cdot 5=300$ models. We fine-tune for a maximum of 20 epochs and early stop when validation loss increases for 3 consecutive epochs.
Figure 8: Sample efficiency of fine-tuning Llama 3 8B on a single task. Hundreds of task examples are typically enough to match or beat the zero-shot performance of GPT-4. Dashed blue line indicates the accuracy of Llama 3 8B fine-tuned on a single task as a function of the number of training examples. The shaded area indicates the 95% confidence interval over the randomly sampled training examples (5 random seeds)
Figure 8 shows how accuracy improves with the number of training examples. Fifty training examples are enough to match or beat the GPT-4 zero-shot baseline for 6 out of the 10 highlighted tasks, and 250 traning examples are enough to match or beat GPT-4 for 8 out of the 10 highlighted tasks. This is crucial, since labeling a few hundred data points is often financially feasible for many legal scholars (Hall and Wright, 2008). With relative few labelled task examples, fine-tuning reasonably small publicly available models can be competitive with state-of-the-art closed models. Moreover, accuracy continues to improve significantly with additional examples. With one thousand training examples, fine-tuning Llama 3 8B Inst matches or beats the GPT-4 baseline for all of the highlighted tasks.

### v2

@@ 2.1 Data sources
2.2 Construction of the classification tasks
What follows is an opinion from the Supreme Court of the United States. Your task
is to identify  whether the opinion effectively says that the decision in this case
"overruled" one or more of the Court\’s own precedents. Alteration also extends to
language in the majority opinion that states that a precedent of the Supreme Court
has been "disapproved," or is "no longer good law". Note, however, that alteration
does not apply to cases in which the Court "distinguishes" a precedent.
@@ 2.2 Construction of the classification tasks
Question: Did the the decision of the court overrule one or more of the Court’s
own precedents?
@@ A. Yes
Think step by step. At the end, respond with "The final answer is [final_answer]",
where [final_answer] is either a single uppercase letter (A-Z) or a numerical value
(e.g., 9, 121).
Figure 3: Example task and prompt corresponding to the Supreme Court “precedent alteration” variable.
We use the variables of the USDB and the USCAD to construct a set of classification tasks. We construct a total of 260 distinct classification tasks, 38 of them corresponding to the Supreme Court database and 232 to the U.S. Court of Appeals. The annotations in the USDB and USCAD serve as labels for these classification tasks. For each task, we construct a prompt template consisting of a general description of the task, followed by a multiple choice question containing each of the possible variable codes. We formulate the task description, question, and answer choices by closely following the language of the variable description in the databases’ documentation. See Figure 3 for an example task, and Appendix D for the full list of tasks.
For every case contained in the SCDB and USCAD, we use the provided case citations to search for its corresponding majority opinion of the court on the Caselaw Access Project , a database of digitized court opinions. Our dataset comprises a total of 24,916 court cases. We divide the court cases into a 70%/10%/20% train/validation/test split. This split enables fine-tuning models for the CaselawQA annotation tasks.
Since many of the classification tasks contain heavily imbalanced classes, we subsample the majority class such that there are at most as many task examples in the majority class as task examples in all other classes combined. As a result, a constant classifier that outputs the majority class label will never achieve more than $50\%$ accuracy on any individual task. This results in a more honest measure of model performance, as models cannot attain high accuracy simply because a task is heavily imbalanced. For completeness, we also report in Appendix B model performance without subsampling of the majority class.
We plot some statistics of the tasks in Figure 2. First, court opinions tend to be long, with 12% having above 8,000 tokens, the typical maximum context size for current state-of-the-art models, such as Llama 3. Second, some tasks have many classes, with 28% of tasks having more than 10 classes. Third, there is a large variability in terms of the number of task examples, ranging from a couple of dozen to 18500 task examples. Our final dataset consists of 718,971 task examples, of which 143,635 belong to the test split.
To reduce the compute required for evaluation, we select at random 5,000 examples from the Supreme Court tasks and 5,000 examples from the Court of Appeals tasks. These 10,000 task examples comprise the test set of CaselawQA, which provides an aggregate measure of model performance. We also make available all 143,635 task examples corresponding to the test court cases, which we denote as the extended test set. Evaluating on the extended test set is 14x as expensive, but provides more fine-grained information on models’ performance across each of the 260 legal classification tasks. We make our classification tasks available as an easy-to-use benchmark within the widely used LM Eval Harness library (Gao et al., 2024).
2.3 Evaluation methodology
We evaluate models using a multiple-choice prompt template similar to the one typically used for the MMLU benchmark (Hendrycks et al., 2020). Since many popular benchmarks are phrased as multiple-choice questions, recent models tend to do well for them (Dominguez-Olmedo et al., 2024). We prompt models to think step-by-step, as we observe that large models achieve higher accuracy when responding with a chain-of-thought (Wei et al., 2022). Due to the diverse set of models and large number of tasks under consideration, we perform no additional task-specific or model-specific prompt tuning.
We use accuracy as the evaluation metric. Since the tasks we consider involve vastly differing numbers of answer choices, accuracy provides an interpretable and comparable measure of performance. Additionally, accuracy is the standard metric used in knowledge-testing LLM benchmarks. For completeness, we also report balanced accuracy and macro-averaged F1 score in Appendix B.
When reporting aggregate performance across multiple tasks (e.g., all Supreme Court tasks), we compute the average accuracy across all task examples. This amounts to a weighted average where tasks are weighted proportional to the number of task examples available.
@@ 2.3 Evaluation methodology
We also highlight throughout the paper ten different annotation tasks that we believe are particularly relevant to the empirical legal research community. Six of these tasks are from the SCDB, and four are from the USCAD. We chose tasks with varying levels of complexity, ranging from relatively simple (e.g., determining the issue area of the case) to more complex (e.g., determining the ideological “direction” of the court decision). See Appendix A for a description of these highlighted tasks.
3 Evaluation baselines
We evaluate the performance of several large language models on CaselawQA, our legal annotation benchmark. Among open-weights models, we evaluate Llama 3.3 70B Instruct (MetaAI, 2024), Qwen 2.5 72B Instruct (Yang et al., 2024), and DeepSeek R1 (Guo et al., 2025), a reasoning model with 685B parameters. For commercial models, we evaluate GPT-4o 2024-08-06 (Hurst et al., 2024), o3-mini 2025-01-31 (OpenAI, 2025b) (high reasoning effort), GPT-4.5 Preview 2025-02-27 (OpenAI, 2025a) and Claude 3.7 Sonnet (Anthropic, 2025) (without reasoning). As a baseline for non-trivial performance, we consider the performance of the constant classifier that always predicts the majority class for each task, regardless of the content of the court opinion being labeled. This simple classifier yields an average accuracy of 40% across the CaselawQA tasks.
Figure 1 shows the accuracy of the evaluated models on CaselawQA. Models with at least 70B parameters achieve accuracies ranging from 70% to 78%, with Claude 3.7 Sonnet performing best. These results indicate that, on average, large models perform well above the constant classifier baseline.
@@ 3 Evaluation baselines
Figure 4: Distribution of task performance across all tasks for Llama 3.1 405B Instruct and Lawma 8B.
However, models display significant variability in performance across legal annotation tasks. Figure 4 (left) shows the distribution of task accuracies for Llama 3.1 405B Instruct—an open-weights model that matches the performance of GPT-4o. Given the high variability in task accuracy, for some of the tasks performance may be sufficient for practical annotation use, though such cases are relatively few. More concerningly, there is a long tail of tasks for which performance is poor. In fact, despite the 405B model achieving an average task accuracy of 76%, it fails to outperform the constant classifier on 58 tasks—22% of the total.
Figure 5: Accuracy of models on the ten highlighted legal classification tasks.
We also examine performance on the subset of 10 highlighted annotation tasks, plotting results for Llama 3.1 405B Instruct and Claude 3.7 Sonnet. Notably, both models show modest performance even on relatively straightforward tasks (e.g.,$<$75 % accuracy on the SC Issue Area task). On more complex tasks, such as SC LC Direction—which involves identifying the ideological direction of a lower court’s decision—models may perform no better than the constant classifier.
Our evaluations indicate that, while large models generally exhibit non-trivial annotation performance, their performance across tasks is highly varied and can be modest even for relatively simple tasks.
Figure 6: For large models, Chain of Thought
prompting outperforms few-shot QA prompting.
Figure 7: Open-weights, general-purpose models for the legal domain perform poorly on CaselawQA.
Few-shot and Chain of Thought prompting.
We now compare Chain of Thought prompring with “Direct QA” prompting, where the model is prompted to directly output an answer label (e.g., “A” or “B”) without producing a reasoning chain. One benefit of Direct QA is that it is straightforward to include examples in-context, whereas few-shot CoT would require collecting reasoning traces for every annotation task.
We compare in Figure 6 the following prompting strategies: zero-shot direct QA, few-shot direct QA, and zero-shot CoT. We consider the Llama 3 Instruct family of models. For these models, we can fit 15 examples in-context, since the model’s maximum context window is 131,072 tokens and each task example is at most 8,192 tokens in length. We observe that for the smaller models (i.e., $\leqslant$3B parameters), few-shot Direct QA performs best. In contrast, for the larger models (i.e., $\geqslant$8B parameters), Chain of Thought prompting is superior. In fact, Llama 3.3 70B Instruct does not benefit from including examples in-context.
Our results indicate that few-shot prompting is not a fruitful strategy to adapt large models to the legal classification tasks at hand. Since court opinions tend to be rather lengthy, few examples may fit in-context, potentially preventing the model from improving over the zero-shot baseline.
Language models for the legal domain.
We additionally evaluate two prominent LLMs adapted to the legal domain: LegalBERT (Chalkidis et al., 2020), a small BERT-style model pre-trained on legal documents, and SaulLM 54B (Colombo et al., 2024a), a Mixtral 7x8B (Jiang et al., 2024) model adapted to the legal domain both through continual pretraining and instruction-tuning on legal text.
We report their CaselawQA performance in Figure 7. We observe that LegalBERT performs poorly, substantially underperforming the constant classifier baseline. This is unsurprising, as LegalBERT is a very small model by today’s standards, with only 110M parameters and a context window of 512 tokens, which most of our Court opinions exceed. Regarding SaulLM 54B, we find that it improves upon its base model–Mixtral 8x7B Instruct– by 4 accuracy points. Nonetheless, its legal annotation performance is poor, and lags that of smaller, generalist models such as Llama 3.1 8B Instruct.
Therefore, existing language models adapted to the legal domain do not perform particularly well on legal annotation tasks, highlighting the need to specialize models for specific annotation tasks of interest.
Figure 8: CaselawQA performance as a function of pre-training compute for Llama 3 Instruct models of varying scales. Performance improves monotonically with compute, but shows steep diminishing returns.
The efficacy of scaling generalist models.
Language models’ downstream benchmark performance generally improves with model scale and pre-training compute (Radford et al., 2019; Brown et al., 2020; MetaAI, 2024). We plot in Figure 8 the CaselawQA performance of the Llama 3 Instruct family of models against their pre-training compute. Similarly to Kaplan et al. (2020), we approximate pre-training compute $C$ in FLOPs as $C\approx 6\cdot N\cdot D$, where $N$ is the model size and $D$ is the number of pretraining tokens.
We find that CaselawQA accuracy improves monotonically with pre-training compute. However, we observe steep diminishing returns. This suggests that further scaling pre-training compute is likely to yield only modest improvements in annotation performance. Additional evidence is that GPT-4.5 performs only slightly better than GPT-4o, see Figure 1, despite being presumably much larger. The limited effectiveness of further scaling generalist models highlights the need to specialize models for legal annotation.
4 The power of specialization
In this section, we present a detailed analysis of how models can be specialized for legal annotation tasks. We start by fine-tuning six different models, ranging in size from 135M parameters to 70B parameters, on all 260 legal annotation tasks from CaselawQA simultaneously, resulting in our Lawma family of models. We then perform additional fine-tuning experiments highlighting different aspects of fine-tuning, such as its sample efficiency, its generalization to unseen tasks and Courts, and the effectiveness of single task specialization.
4.1 The Lawma models
Figure 9: Performance of the Lawma models. The smallest Lawma model, Lawma 135M, outperforms both largest Llama 3 model, Llama 3.1 405B, and the best-performing commercial model, Claude 3.7 Sonnet.
We first fine-tune on all CaselawQA tasks simultaneously. We fine-tune the following models: HuggingFace’s SmolLM2 135M and 360M Instruct (Allal et al., 2025), Llama 3.2 1B and 3B Instruct, Llama 3.1 8B Instruct, and Llama 3.3 70B Instruct (MetaAI, 2024). We refer to these models as Lawma 135M, Lawma 360M, Lawma 1B, Lawma 3B, Lawma 8B, and Lawma 70B respectively. The fine-tuning dataset contains approximately 553,000 tasks examples, totalling 2.2B tokens. We fine-tune for 3 epochs. We find that additional epochs do not significantly improve performance. See Appendix C for additional details regarding the model training.
We compare in Figure 9 the performance of the Lawma models on CaselawQA with Llama 3.1 405B Instruct—the largest Llama 3 model—and Claude 3.7 Sonnet, the best-performing commercial model. The largest Lawma model, Lawma 70B, achieves 88% accuracy, outperforming all commercial models by at least 10 percentage points. Remarkably, even the smallest Lawma model, Lawma 135M, reaches 83% accuracy—surpassing all commercial models despite being over three orders of magnitude smaller.
The performance of the Lawma models improves consistently with model size. However, these gains are relatively modest: tripling model size generally yields an accuracy increase of only about 1%. Notably, the Lawma 8B model achieves 86% accuracy while fitting on a single H100 GPU, offering a favorable trade-off between performance and efficiency. We therefore focus on Lawma 8B in the remainder of our analysis.
For the 10 legal annotation tasks highlighted in Figure 5, Lawma 8B outperforms both Claude 3.7 Sonnet and Llama 3.1 405B Instruct on 7 tasks and matches their performance on the remaining 3. In some cases, the improvements are remarkably large: for instance, Lawma 8B outperforms Claude 3.7 Sonnet by 30 percentage points on the SC LC Direction task and by 20 points on Songer Case Origin.
Finally, Figure 4 (right) shows the distribution of task accuracies for Lawma 8B. Most task accuracies lie between 75% and 100%, suggesting that, compared to commercial models, Lawma 8B may be suitable for practical annotation use across a substantially wider range of tasks. Nonetheless, accuracy varies considerably across tasks, and the model performs poorly for a sizable number of annotation tasks.
@@ 4.2 Sample efficiency
Figure 10: Sample efficiency of fine-tuning Llama 3.1 8B Instruct on a single task. Hundreds of task examples can be enough to match or beat the performance of Claude 3.7 Sonnet. The dashed blue line indicates the accuracy of the fine-tuned model as a function of the number of training examples. The shaded area indicates the 95% confidence interval over the randomly sampled training examples (5 random seeds)
We investigate how task accuracy scales as models are trained on increasing numbers of examples. Specifically, we focus on the 10 highlighted annotation tasks. We independently fine-tune the Llama 3.1 8B Instruct model for each task, in contrast to previous experiments where models were trained on all tasks simultaneously.
We fine-tune on 10, 50, 100, 250, 500, and 1000 examples per task, selecting examples uniformly at random. For each, we train 5 models using different random seeds. Each model is trained for up to 20 epochs, with early stopping triggered if the validation loss increases for three consecutive epochs.
Figure 10 shows how accuracy improves with more training data. For comparison, we include the performance of Claude 3.7 Sonnet. Remarkably, with 1000 or fewer labeled examples, our fine-tuned models match or exceed Claude 3.7 Sonnet on 7 of the 10 highlighted tasks. This is crucial, since labeling hundreds of data points is often financially feasible for many legal scholars (Hall and Wright, 2008). With relatively little labeled data, fine-tuning small open-weights models can offer competitive performance relative to much larger models. Furthermore, accuracy continues to improve with additional examples.
4.3 Generalization to unseen legal databases
Figure 11: Fine-tuning Llama 3.1 8B IT only on Court of Appeals tasks improves performance on Supreme Court tasks. However, it still lags significantly behind Claude 3.7 Sonnet and Lawma 8B.
Figure 12: The performance of Lawma 8B trails behind inter-coder agreement for many Court of Appeals tasks.
We now examine the extent to which fine-tuning generalizes to unseen legal databases. Specifically, we fine-tune Llama 3.1 8B Instruct for one epoch on all Court of Appeals (CoA) tasks simultaneously, then evaluate its average performance on all Supreme Court (SC) tasks. As baselines, we use the original Llama 3.1 8B Instruct and Claude 3.7 Sonnet—neither of which were fine-tuned—as well as Lawma 8B, which was fine-tuned on all annotation tasks, including those from the Supreme Court database.
Figure 11 shows the average accuracy on the Supreme Court tasks. Fine-tuning on CoA tasks alone leads to a notable improvement, raising accuracy from 50% to 64%. This result indicates that specializing models for legal annotation can improve performance on unseen annotation tasks and databases.
However, the model fine-tuned only on CoA tasks still underperforms compared to Claude 3.7 Sonnet by 16 percentage points. Likewise, Lawma 8B, which was fine-tuned on SC tasks, outperforms the CoA-only fine-tuned model by 25 percentage points. These findings underscore the importance of fine-tuning models on the precise target annotation tasks of interest. “Broadly specializing” a model for legal annotation is not sufficient; targeted fine-tuning is essential. This also suggests that the Lawma models may not outperform commercial models on annotation tasks beyond those in the Court of Appeals and Supreme Court databases.